OpenStudy (anonymous):

Solve this system of equations: 2x-5y+4z=18 x+3y+5z=7 -3x-2y+4z=22

OpenStudy (anonymous):

@ganeshie8 im bothering u alot today, but help please >.<

ganeshie8 (ganeshie8):

hey np, gimme one sec

OpenStudy (anonymous):

ok thank u :)

ganeshie8 (ganeshie8):

\[ 2x-5y+4z=18 -------------- (1)\\ x+3y+5z=7 -----------------(2)\\ -3x-2y+4z=22 ------------(3) \\ \]

OpenStudy (anonymous):

is that all i have to put?

ganeshie8 (ganeshie8):

lol you're kidding... we need to solve x, y, z here

ganeshie8 (ganeshie8):

we do that step by step, do you need to show work. or is this a multiple choice q ?

OpenStudy (anonymous):

its not multiple choice but i dont needa show work either.

ganeshie8 (ganeshie8):

then you can use wolfram

OpenStudy (anonymous):

oh okay true how do i plug it in?

ganeshie8 (ganeshie8):

separate the equations with commas and plug in. like this :- http://www.wolframalpha.com/input/?i=2x-5y%2B4z%3D18%2C++++x%2B3y%2B5z%3D7%2C++++-3x-2y%2B4z%3D22

ganeshie8 (ganeshie8):

wolfram is super smart, since you're graduating in a while its good you get a hang of it for problems like these. you can avoid the donkey work..

OpenStudy (anonymous):

cool thanks for the website & help! i got it.

ganeshie8 (ganeshie8):

great !

ganeshie8 (ganeshie8):

but use it, only when you are familiar wid the method of solving these equations by yourself. that is, use it when you feel lazy ;p

OpenStudy (anonymous):

haha okaay i sure will :p

ganeshie8 (ganeshie8):

i trust you i hope i didnt do a mistake by guiding you to that website :S

OpenStudy (anonymous):

no i also know another website called mathway that solves problems for you too but i only use it when neccesary :p so ur trust is worthy in me! lol

ganeshie8 (ganeshie8):

sounds we're good :)

ganeshie8 (ganeshie8):

and just to show u how smart wolfram is :- you posted a question asking for a line with slope -3/4, and goes through point (5,9) : see, what wolfram gives :- http://www.wolframalpha.com/input/?i=slope+-3%2F4%2C+point+%284%2C+5%29

ganeshie8 (ganeshie8):

* point (4, 5)

OpenStudy (anonymous):

Ooo ok i see but do i put the equation form that wolfram is showing or the other person that replied? :o

ganeshie8 (ganeshie8):

ok ima answer it here

ganeshie8 (ganeshie8):

the q is asking you for standard form. do you knw what a standard form is ?

OpenStudy (anonymous):

yes like typed out/written out.

ganeshie8 (ganeshie8):

standard form of a line is this : Ax + By = C

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

3x+4y-32 = 0 3x + 4y = 32

ganeshie8 (ganeshie8):

thats the standard form of line

OpenStudy (anonymous):

so those two are my answrs

ganeshie8 (ganeshie8):

only one is our answer :

ganeshie8 (ganeshie8):

3x + 4y = 32

ganeshie8 (ganeshie8):

Standard Form must have the form : Ax + By = C

OpenStudy (anonymous):

3x+4y=32?

OpenStudy (anonymous):

oh ok got it!

ganeshie8 (ganeshie8):

yes ! finally ;)

OpenStudy (anonymous):

:D